OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. Choose one answer. a. y - 2 = 3(x - 5) b. y + 8 = 3(x + 5) c. y - 2 = 3(x - 1) d. y - 5 = 3(x - 2)

OpenStudy (saifoo.khan):

how many do u have? :O

OpenStudy (anonymous):

a few more >.< but thank you so much for all the help you have given me!

OpenStudy (saifoo.khan):

uw, i can help u.. but they are a bit lengthy! And i hate them, lol.

OpenStudy (anonymous):

yes they are very lengthy and i tend to get soo flustered with them....i have problems with numbers.....

OpenStudy (saifoo.khan):

nvm, i will solve!

OpenStudy (saifoo.khan):

:]

OpenStudy (saifoo.khan):

d. y - 5 = 3(x - 2)

OpenStudy (saifoo.khan):

ans.

OpenStudy (zarkon):

point slope form of a line \[y-y_1=m(x-x_1)\] \[m=3,x_1=2,y_1=5\] has to be (d)

OpenStudy (anonymous):

thank you!!!

OpenStudy (saifoo.khan):

uw!!!